Please help me solve this equation:
3x-4y=2
2x-4y=-4

Answer by mananth(16946)

3 x -4 y = 2 .............1
Total value
2 x -4 y = -4 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-3 x 4 y = -2
2 x -4 y = -4
Add the two equations
-1 x = -6
/ -1
x = 6
plug value of x in (1)
3 x -4 y = 2
18 -4 y = 2
-4 y = 2 -18
-4 y = -16
y = 4